This is a projectile motion question from my physics test I need help with for when I make test corrections. I only have 40 minutes for test corrections later today, so I could only memorize the question. Please bear with me here.

"A skier goes off a jump that is 2.5 meters long, and the jump is at a 30 degree angle. The initial velocity of the skier is 16.2m/s. Assume wind resistance is negligible. What is the height from the top of the jump to the ground and at what point will the skier hit the ground?"

I'm pretty sure the height of the jump will be 1.25 meters but am not sure how to find the second part of the answer.

Please show your work.
Answer #1

sx = 2.5 m
angle = 30 degrees
u = 16.2
u = 16.2cos30 i + 16.2 sin30 j
u = 14.096 i + 8.1 j

Height of the jump = 2.5*sin30 = 1.25

Now
while climbimg the height 1.25 m the veritcal velocity will reduce.
Let the launch velocity be ly
v^2 = u^2 + 2as
ly^2 = 8.1^2 - 2*9.8*1.25
ly^2 = 41.11
ly = 6.4117 m/s

Time taken to reach the maximum height =>
v = u + at
0 = ly - 9.8*t
t = 6.4117/9.8
t = 0.654 sec

Hence, the horiziontal distance travelled :
sx = 14.096*0.654 = 9.22 m

Please give a thumps up if the answer is correct.